How many pounds of extra-lean hamburger that is 7%fat must be mixed with hamburger that is 15% fat to obtain a mixture of 30 pounds of hamburger that is 10% fat?
Answer by Boreal(15235)

x at 7%
30-x at 15%
.07x+.15(30-x)=30*0.10=3 pounds pure fat
.07x+4.5-0.15x=3
-0.08x=-1.5
x=18.75 pounds of 7% (which is 1.3125 lb pure fat)]
30-x=11.25 pounds of 15% (which is 1.6875 lb pure fat)